OpenStudy (anonymous):

Find exact value if: 0 < x < pi/2. and 0 < y < pi/2 Sin ( x-y) if sin x = 3/5 and sin y= 4/5

OpenStudy (anonymous):

sin(x-y)=sin(x)cos(y)-cos(x)cos(y) sin(x)=3/5 ==> cos(x)=4/5, and sin(y)=4/5 ==> cos(y)=3/5 sin(x-y)=(3/5)(4/5)-(4/5)((3/5) =0

OpenStudy (anonymous):

your answer is the same my,but the key is :24/25

OpenStudy (anonymous):

that would be the answer for sin(x+y)

OpenStudy (anonymous):

tk any way

OpenStudy (anonymous):

I check (x-y) not (x+y) don't worry

OpenStudy (anonymous):

tk any way

OpenStudy (anonymous):

your answer is the same my,but the key is :24/25